OpenStudy (anonymous):

write the algebraic equation Seven less than a number is 15.

sam (.sam.):

-8

OpenStudy (anonymous):

if u want me to answer give me a medal first!!!!! kkk and i also gave u a medal for ur question cuz im so kind.

OpenStudy (anonymous):

the answer is

OpenStudy (anonymous):

x-7=15 x=22

sam (.sam.):

7-(x)=15 -x=8 x=-8

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

it cant be -8. the seven less than the number came first..its 22. (x) is the "number"